Systems are not running as smoothly as I would like! While most of AAVSOnet is producing good data, here are a few instances where you need to watch out.
BSM_Hamren. Down until the mount gets fixed.
TMO61. We replaced the QSI-683 camera with the older QSI-583. The hopes are that this will solve the camera connect issues we have been having. We have also added a LISA spectrograph that can be swapped in for bright-time. However, now the monsoon has hit with full force. Don't expect much data from TMO61 until September.
BSM_NM. The summer monsoon has arrived. Don't expect much data until September.
SRO. The summer monsoon has arrived. Don't expect much data until September.
OC61. For the past two weeks, there has been a drive issue. This telescope has a "belt drive", using a stainless steel drive belt over smooth wheels. In general, this gives reasonable drive performance without any periodic error. However, one of the bearings had become stiff and was causing the belt to slip. On some nights, we were pointing many degrees off from the correct position. This problem has now been identified and hopefully solved. At the same time, the weather has been abysmal, so there really hasn't been much of a loss of actual telescope time.
BSM_South and BSM_Berry. These southern systems have been used for the past 3 weeks to perform time series on a setting object. Any program target in the first half of the night has therefore been missed. We'll take a break from this starting tomorrow and let some of the monitoring projects have priority.
coker30. This system was down for about 3 weeks between 5/30 and 6/18, primarily so the telescope operator could attend meetings. It has been back in full operation since 6/18, and will be our main SW US system until the AZ/NM summer monsoon ends.